IF THE TOTAL COST FUNCTION OF A FIRM IS GIVEN AS TC= Q raise to power 3+8Q raise to power 2+ 6Q+200


Determine the function of 1. AVERAGE VARIABLE COST

2.AVERAGE TOTAL COST

3.AVERAGE FIXED COST

4. MARGINAL COST @ Q=5


1
Expert's answer
2020-12-11T11:09:44-0500

TC = Q3+8Q2+6Q+200

Average Total cost = TC / Q = Q2+8Q+6+200/Q

Average Fixed cost = 200/Q

Average Variable cost = Q2+8Q+6

MC = d(TC)/dQ = 3Q2+16Q+6

Q @ 5 , MC = 3(5)2+15(5)+6 = 156